Assume that a real estate manager has 5 contacts per week, and she believes that for each contact the probability of making a sale is 0.30. Calculate the following if possible. If any part of question

(you feel) is not possible to calculate give reason:

a. Develop the probability distribution table for sale

b. Calculate the expected number of sales per week and variance of sale

c. Find the probability that she makes at most 2 sales.

d. Find the probability that she makes between 3 and 6 sales (both inclusive)


1
Expert's answer
2020-11-24T16:15:16-0500

a)



b)"\mu" (expected number of sales per week=0.36015+0.3087*2+0.1323*3+0.02835*4+0.00243*5)=1.5

variance = "\sum" pi(xi-"\mu")2 = 0.16807*(0-1.5)2+0.36015*(1-1.5)2+0.3087*(2-1.5)2+0.1323*(3-1.5)2+0.02835*(4-1.5)2+0.00243*(5-1.5)2 = 1.05

c)0.16807+0.36015+0.3087=0.83692

d)0.1323+0.02835+0.00243=0.16308